### 13 Dec 2023 v1

# I Introduction

Abstract
LiGaO in the spinel type structure is investigated as a potential ultra-wide-band-gap semiconductor. The band structure is determined using the quasiparticle self-consistent method and the optical properties are calculated at the Bethe Salpeter Equation level including electron-hole interaction effects. The optical gap including exciton effects and an estimate of the zero-point motion electron phonon coupling renormalizations is estimated to be about 5.2 eV with an exciton binding energy of about 0.4 eV. Si doping as potential n-type dopant is investigated and found to be a promising shallow donor.
I Introduction
There is a need for ultra-wide-band-gap semiconductors both for optoelectronic and high-power applications. The high-power applications benefit from a large gap because it is found that the breakdown fields scale with the band gap. Recently, there has been a lot of interest in -GaO, because of its gap of 4.9 eV.[1, 2] In spite of such a large gap it can be doped with Si, Ge or Sn to give semiconducting properties. This finding has sparked a world-wide development of this material in epitaxial form and heterojunction devices involving even higher gap (AlGa)O are making rapid progress. However, -type doping remains very difficult of elusive, related to self-trapped hole polaron formation, related in turn to high valence band effective masses. It has nonetheless sparked interest in fundamental properties because of the unusual monoclinic structure with both tetrahedral and octahedral Ga sites.
On the other hand, LiGaO (lithium gallate) has an even wider band gap and was recently also predicted to be -type dopable by Si or Ge. [3, 4] Like GaO bulk crystals can be grown and offer the possibility of homo-epitaxial growth. It has a much simpler wurtzite based crystal structure with all Ga and Li in tetrahedral coordinations and can be thought of as derived from ZnO by replacing the group-II ion Zn by alternating group-I (Li) and group III Ga ions. Further band gap tuning is possible by replacing Ga by Al and Li by Na in the same crystal structure or closely related structures.[5, 6] The anisotropy of the material leads to interesting exciton splittings with large exciton binding energies and recent calculations [7] were found to be in excellent agreement with experiment[8, 9, 10] and give an exciton gap of 6 eV. It thus appears useful to develop epitaxial growth and doping of this material to pursue it as an active semiconductor material and not merely an optical material.
In the Li-Ga-O ternary system, another stable compound exists with formula LiGaO. One can think of it as taking 4 LiGaO units and replacing 3 Li with one Ga to maintain the charge balance. However, LiGaO is known to have a different crystal structure. It has a cubic spinel-type structure with Ga occurring in both tetrahedral and octahedral sites and Li in octahedral sites.[11] Its structure and related information can be found in [12, 13]. We used the experimental crystal structure from the International Crystallographic Database (ICSD),[14] as shown in Fig. 1. Interestingly, a defective spinel phase also occurs in GaO as a metastable phase and has been labeled the -phase [15]. It has recently been observed to occur during certain processing steps in doped -GaO and a model for how the transition from to phase can proceed via a succession of defect formations was proposed by Hsien-Lien et al. [16, 17] Since LiGaO is a compound with composition in between GaO and LiGaO it is of interest to determine its band structure and optical properties to assess its potential as UWBG semiconductor. It has previously been studied as a phosphor when doped with Cr.[18, 19, 20] Very recently, LiGaO was grown epitaxially and reported to exchibit -type conduction.[21] The relation to the -phase of GaO, which is known to have a higher band gap than the -phase, provides additional interest. It suggests that adding a small amount of Li may stabilize the -phase.
This paper presents results on the electronic band structure using the state-of-the-art quasiparticle self-consistent method and also calculates optical properties, including excitonic effects. Since the unit cell already has 56 atoms, (8 formula units) we can already obtain an idea of doping by simply replacing one Ga by a Si. We thus also study a LiGaSiO cell with Si placed on either a tetrahedral or octahedral Ga site.
II Computational Methods
The calculations were performed using the Questaal suite,[22] which use the full-potential linearized muffin-tin orbital method(FP-LMTO) [23] to implement both density functional theory (DFT) and many-body perturbation theory (MBPT) approaches such as Hedin’s method for quasiparticle band structures ( for Green’s function and screened Coulomb interaction)[24, 25] in a quasiparticle self-consistent version called QS, [26, 27] and the Bethe Salpeter Equation (BSE) approach for optical response calculations.[28, 29, 30] Specifically, we here use an extension of the QS method in which the screened Coulomb interaction is calculated including ladder diagrams, which has been dubbed QS,[30] and involves solving a BSE equation not just in the limit for optical response but for the mesh of for which is needed. The basis set angular momentum cut-offs for first and second set of smoothed Hankel function basis sets were =3,3 for Ga and O and 3,2 for Li. Augmentations inside the spheres were carried out till . In the QS calculations 32 valence bands and 4 conduction bands were included in the active space in which the ladder diagrams are evaluated. The PBEsol functional was used for the initial DFT calculations.[31]. The up to which self-energies were carried out was set to 3.5 Ry and the self-energy matrix was approximated by its diagonal above 3.0 Ry with an average value used evaluated over the range Ry.
III Results
The band structure of LiGaO is shown in Fig. 2 as obtained in the GGA and QS methods. The conduction band minimum (CBM) occurs at and the top valence band is very flat but with valence band maximum (VBM) at as shown in Fig. 3. The gap is thus slightly indirect with indirect gap of 5.72 eV and direct gap at at 5.84 eV. The gap is severely underestimated by GGA. Some further detail on the gap are given in Table 1. The gap shift from GGA to QS is 3.24 eV and results from a downward shift of the VBM by 1.43 eV and 1.81 upward shift of the CBM.
There is here barely any difference between QS and QS but this may be a result of including insufficient number of bands in the active space. For LiGaO, we also found a rather small reduction of the gap correction of about 0.95 compared to usual 0.8. Here it is only 0.99. This may be related to the number of bands included in the active space for which the BSE two-particle Hamiltonian is diagonalized. Unfortunately, including all O- bands would require 96 valence bands and including at least one band per Ga in the conduction band would require 20 conduction bands. This becomes rather challenging in terms of memory allocation. To test the idea, we repeated some of our previous calculations of LiGaO but with a higher number of conduction bands. Instead of using all 52 occupied bands (20 Ga- like, 8 O- and 24 O-) and only 4 conduction bands which was the default used in [7] we now use 24 valence bands and 12 conduction bands. This gives a gap of 6.737 eV instead of 7.016 eV in [7]. With a QSGW gap of 7.218 eV and GGA gap of 3.314 eV, the reduction factor or rather of 0.88 where we use the estimate for either or . While this reduces the QP-gap of LiGaO by about 0.3 eV, the same reduction of also affects the exciton binding energy. Using a similar procedure as in [7] we then find an extrapolated exciton gap of 6.4 eV which with a zero-point motion correction still gives a lowest exciton gap of 6.0 eV in good agreement with the experimental data.[10] Nonetheless, it would mean that the exciton binding energy is only about 0.3 eV instead of 0.7 eV. The qualitative results of Ref. [7] about the existence of a Rydberg series of excitons stay valid. This suggests that we should apply a somewhat larger reduction factor of the gap correction than found in Table 1. Using a reduction factor due to ladders of 0.88 we would obtain a indirect gap of 5.4 eV and direct gap at of 5.5 eV.
| GGA | QS | QS | |
| indirect () | 2.48 | 5.75 | 5.72 |
| direct | 2.58 | 5.87 | 5.84 |
The dielectric function is shown in Fig. 4 in the independent particle approximation (IPA) and the BSE approximation. We can see a significant effect from including the electron-hole interaction effects and excitons with a large exciton binding energy. We caution that the dielectric constant obtained here is likely overestimated as is the intensity of because of difficulties in calculating the contribution of the self-energy to the velocity matrix elements but the peak positions should be accurate. In this calculation we only used in a mesh for the BSE equation. To obtain the excitons more accurately, we calculated the lowest exciton for finer k-meshes with but with smaller number of bands 10 valence bands instead of 32 while keeping 4 conduction bands. We can see that reducing the number of valence bands makes a difference of only 0.01 eV with slightly lower gaps obtained with more valence bands included. Extrapolating the gap as function of , i.e. the total number of mesh points in the BZ, as we did in [7] we find a extrapolated lowest exciton at 5.48 eV, close to our calculated gap for . This would amount to a exciton binding energy of 0.4 eV. The BSE calculations here only include direct transitions. Therefore we deduce the exciton binding energy with respect to the direct gap at . One may expect a similar binding energy for an exciton related to the indirect gap which is 0.1 eV smaller. Neither the quasiparticle gap nor the exciton gap here include electron-phonon coupling renormalization. In LiGaO, the latter was found to be eV. A similar value is expected here and would reduce our direct gap of Table 1 from 5.87 to 5.60.1 eV (quasiparticle ) and 5.2.0.1 eV (exciton) and another 0.1 eV lower for the indirect quasiparticle and exciton gaps. Both the quasiparticle gap and exciton binding energy could be somewhat lower if we assume a larger reduction of relative to due to ladder diagrams but as mentioned earlier for LiGaO, we expect these two errors to compensate each other and give a similar optical gap. This gap is slightly higher than the gap of 5.0 eV reported [15] for -GaO.
To evaluate the n-type dopability of the system, we replaced one Ga atom in the 56 atom cell with Si. Fig.6(a) shows the band structure for the tetrahedral Si site after relaxation and in the GGA. One can see that the extra valence electron essentially starts filling the conduction band but no levels occur in the gap. The gap in fact slightly increased from 2.58 eV in pure LiGaO in the GGA to 2.74 eV. This is partly due to the relaxation. Without relaxation the gap would have decreased slightly to 2.39 eV. The Si contribution to the bands in this figure is indicate by the red color while the background color of the bands without Si contribution is grey. The relaxation is significant. As can be seen in Fig. 7, the nearest neighbor O move inward toward the Si but with also a distortion and symmetry lowering making one Si-O bond longer than the other three. The bond lengths change from 1.776 Å for the Ga-O tetrahedral bond to 1.675 Å (3 bonds) and 1.695 Å (one Si-O) bond. The Ga atoms bonded to these nearest neighbor O atoms then have a larger bond length again by about 0.1 Å than the original octahedral Ga-O bond length of 2.05 Å. In spite of these relaxations there clearly is no deep level formation. This indicates that Si would be a shallow donor and could lead to n-type doping.
We have also calculated Si on an octahedral site with similar results. First, we find that the relaxed total energies were within error bar indistinguishable. So Si is likely to occupy either octahedral or tetrahedral sites with equal likelihood. Secondly, we also find no states in the gap indicating Si is a shallow donor on both sites as seen in
Fig. 6(b). Similar to the tetrahedral case, we find a slight increase in gap due to the Si replacement to 2.688 eV in GGA. Two of the octahedral bonds around Si are 1.814 Å, two are 1.864 Å, and two are 1.784Å. The Ga-O octahedral bonds in the host are 1.982 Å. Thus, there is again an inward relaxation of nearby O toward the Si.
While we have here only studied the Si doping via the GGA band structure, we expect that the shallow donor character will stay valid at the QS level. A similar situation was found recently for LiAlO.[5]
IV Conclusion
In summary, LiGaO in the spinel type cubic structure is a good candidate UWBG semiconductor. It is predicted to have slightly higher gap than -GaO and but not as high as LiGaO and appears to be n-type dopable by Si on either its tetrahedral or octahedral Ga site. Its cubic structure may have advantages.
Acknowledgements.
This work made use of the High Performance Computing Resource in the Core Facility for Advanced Research Computing at Case Western Reserve University and was supported by the U.S. Air Force Office of Scientific Research (AFOSR) under grant no. FA9550-22-1-0201. Data Availability: The data that supports the findings of this study are available within the articleReferences
- Matsumoto et al. [1974] T. Matsumoto, M. Aoki, A. Kinoshita, and T. Aono, Absorption and Reflection of Vapor Grown Single Crystal Platelets of -Ga 2 O 3, Jap. J. Appl. Phys. 13, 1578 (1974).
- Sasaki et al. [2013] K. Sasaki, M. Higashiwaki, A. Kuramata, T. Masui, and S. Yamakoshi, ”{MBE} grown ga2o3 and its power device applications ”, J. Cryst. Growth 378, 591 (2013), the 17th International Conference on Molecular Beam Epitaxy.
- Boonchun et al. [2019] A. Boonchun, K. Dabsamut, and W. R. L. Lambrecht, First-principles study of point defects in LiGaO, Journal of Applied Physics 126, 155703 (2019).
- Dabsamut et al. [2020] K. Dabsamut, A. Boonchun, and W. R. L. Lambrecht, First-principles study of n- and p-type doping opportunities in LiGaO, Journal of Physics D: Applied Physics (2020).
- Popp and Lambrecht [2022] P. Popp and W. R. L. Lambrecht, Quasiparticle self-consistent band structures and phase transitions of in tetrahedrally and octahedrally coordinated structures, Phys. Rev. Mater. 6, 104605 (2022).
- Radha et al. [2021] S. K. Radha, A. Ratnaparkhe, and W. R. L. Lambrecht, Quasiparticle self-consistent band structures and high-pressure phase transitions of and , Phys. Rev. B 103, 045201 (2021).
- Dadkhah et al. [2023] N. Dadkhah, W. R. L. Lambrecht, D. Pashov, and M. van Schilfgaarde, Improved quasiparticle self-consistent electronic band structure and excitons in , Phys. Rev. B 107, 165201 (2023).
- Trinkler et al. [2017] L. Trinkler, A. Trukhin, B. Berzina, V. Korsaks, P. Ščajev, R. Nedzinskas, S. Tumėnas, M. Chou, L. Chang, and C.-A. Li, Luminescence properties of LiGaO crystal, Optical Materials 69, 449 (2017).
- Trinkler et al. [2022] L. Trinkler, V. Pankratov, A. Trukhin, B. Berzina, M. Chou, and L. Chang, Anisotropic photoluminescence of -LiGaO crystal, Optical Materials 132, 112856 (2022).
- Tuménas et al. [2017] S. Tuménas, P. Mackonis, R. Nedzinskas, L. Trinkler, B. Berzina, V. Korsaks, L. Chang, and M. Chou, Optical properties of lithium gallium oxide, Applied Surface Science 421, 837 (2017).
- Joubert et al. [1963] J. Joubert, M. Brunel, A. Waintal, and A. Durif, Etude cristallographique du gallate de lithium et de sa solution solide avec l‘aluminate, Comptes Rendus Hebdomadaires des Seances de l’Academie des Sciences (1884 - 1965) 256, 5324–5326 (1963).
- [12] Data retrieved from Materials Project for LiGa5O8 (mp-28146) from database version v2022.10.28.
- Jain et al. [2013] A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder, and K. A. Persson, Commentary: The Materials Project: A materials genome approach to accelerating materials innovation, APL Materials 1, 011002 (2013).
- [14] International Crystallographic Data Base, id=33716.
- Biswas and Nishinaka [2022] M. Biswas and H. Nishinaka, Thermodynamically metastable -, - (or -), and -Ga2O3: From material growth to device applications, APL Materials 10, 060701 (2022), https://pubs.aip.org/aip/apm/article-pdf/doi/10.1063/5.0085360/16490464/060701_1_online.pdf .
- Huang et al. [2023a] H.-L. Huang, C. Chae, J. M. Johnson, A. Senckowski, S. Sharma, U. Singisetti, M. H. Wong, and J. Hwang, Atomic scale defect formation and phase transformation in Si implanted -Ga2O3, APL Materials 11, 061113 (2023a), https://pubs.aip.org/aip/apm/article-pdf/doi/10.1063/5.0134467/17941303/061113_1_5.0134467.pdf .
- Huang et al. [2023b] H.-L. Huang, J. M. Johnson, C. Chae, A. Senckowski, M. H. Wong, and J. Hwang, Atomic scale mechanism of to phase transformation in gallium oxide, Applied Physics Letters 122, 251602 (2023b), https://pubs.aip.org/aip/apl/article-pdf/doi/10.1063/5.0156009/18006597/251602_1_5.0156009.pdf .
- De Clercq et al. [2017] O. Q. De Clercq, L. I. D. J. Martin, K. Korthout, J. Kusakovskij, H. Vrielinck, and D. Poelman, Probing the local structure of the near-infrared emitting persistent phosphor LiGa5O8:Cr3+, J. Mater. Chem. C 5, 10861 (2017).
- Sousa and Carvalho [2020] O. Sousa and I. Carvalho, Theoretical study of the structural, energetic, electronic and magnetic properties of the host matrix LiGa5O8 doped with Cr3+, Journal of Solid State Chemistry 289, 121472 (2020).
- Huang et al. [2018] W. Huang, X. Gong, R. Cui, X. Li, L. Li, X. Wang, and C. Deng, Enhanced persistent luminescence of LiGa5O8:Cr3+ near-infrared phosphors by codoping Sn4+, Journal of Materials Science: Materials in Electronics 29, 10535 (2018).
- [21] K. Zhang, V. G. T. Vangipuram, H.-L. Huang, J. Hwang, and H. Zhao, Discovery of a Robust P-Type Ultrawide Bandgap Oxide Semiconductor: LiGa5O8, Advanced Electronic Materials n/a, 2300550, https://onlinelibrary.wiley.com/doi/pdf/10.1002/aelm.202300550 .
- Pashov et al. [2019] D. Pashov, S. Acharya, W. R. Lambrecht, J. Jackson, K. D. Belashchenko, A. Chantis, F. Jamet, and M. van Schilfgaarde, Questaal: A package of electronic structure methods based on the linear muffin-tin orbital technique, Computer Physics Communications , 107065 (2019).
- Kotani and van Schilfgaarde [2010] T. Kotani and M. van Schilfgaarde, Fusion of the LAPW and LMTO methods: The augmented plane wave plus muffin-tin orbital method, Phys. Rev. B 81, 125117 (2010).
- Hedin [1965] L. Hedin, New method for calculating the one-particle green’s function with application to the electron-gas problem, Phys. Rev. 139, A796 (1965).
- Hedin and Lundqvist [1969] L. Hedin and S. Lundqvist, Effects of electron-electron and electron-phonon interactions on the one-electron states of solids, in Solid State Physics, Advanced in Research and Applications, Vol. 23, edited by F. Seitz, D. Turnbull, and H. Ehrenreich (Academic Press, New York, 1969) pp. 1–181.
- van Schilfgaarde et al. [2006] M. van Schilfgaarde, T. Kotani, and S. Faleev, Quasiparticle Self-Consistent Theory, Phys. Rev. Lett. 96, 226402 (2006).
- Kotani et al. [2007] T. Kotani, M. van Schilfgaarde, and S. V. Faleev, Quasiparticle self-consistent GW method: A basis for the independent-particle approximation, Phys.Rev. B 76, 165106 (2007).
- Onida et al. [2002] G. Onida, L. Reining, and A. Rubio, Electronic excitations: density-functional versus many-body Green’s-function approaches, Rev. Mod. Phys. 74, 601 (2002).
- Cunningham et al. [2018] B. Cunningham, M. Grüning, P. Azarhoosh, D. Pashov, and M. van Schilfgaarde, Effect of ladder diagrams on optical absorption spectra in a quasiparticle self-consistent framework, Phys. Rev. Materials 2, 034603 (2018).
- Cunningham et al. [2023] B. Cunningham, M. Gruening, D. Pashov, and M. van Schilfgaarde, QS: Quasiparticle Self consistent GW with ladder diagrams in W (2023), arXiv:2302.06325 [cond-mat.mtrl-sci] .
- Perdew et al. [2008] J. P. Perdew, A. Ruzsinszky, G. I. Csonka, O. A. Vydrov, G. E. Scuseria, L. A. Constantin, X. Zhou, and K. Burke, Restoring the Density-Gradient Expansion for Exchange in Solids and Surfaces, Phys. Rev. Lett. 100, 136406 (2008).